# Matryoshka Representations for Adaptive Deployment

### v2

@@ 2 Related Work
$\rm MRL$} can adapted seamlessly to most representation learning frameworks at web-scale with minimal modifications (Section 4.1). For example, $\rm MRL$}’s adaptation to masked language modelling reduces to $\rm MRL\text{--}E$} due to the weight-tying between the input embedding matrix and the linear classifier. For contrastive learning, both in context of vision & vision + language, $\rm MRL$} is applied to both the embeddings that are being contrasted with each other. The presence of normalization on the representation needs to be handled independently for each of the nesting dimension for best results (see Appendix C for more details).
@@ 4.2.1 Adaptive Classification
Figure 7 shows the comparison between cascaded $\rm MRL$} representations ($\rm MRL$}–AC) and independently trained fixed feature (FF) models on ImageNet-1K with ResNet50. We compute the expected representation size for $\rm MRL$}–AC based on the final dimensionality used in the cascade. We observe that $\rm MRL$}–AC is as accurate, $76.30\%$, as a 512-dimensional FF model but requires an expected dimensionality of $\sim 37$ while being only $0.8\%$ lower than the 2048-dimensional FF baseline. Note that all $\rm MRL$}–AC models are significantly more accurate than the FF baselines at comparable representation sizes. $\rm MRL$}–AC uses up to $\sim 14\times$ smaller representation size for the same accuracy which affords computational efficiency as the label space grows [80]. Lastly, our results with $\rm MRL$}–AC indicate that instances and classes vary in difficulty which we analyze in Section 5 and Appendix J.
@@ 4.3 Retrieval
$\rm MRL$} models are capable of performing accurate retrieval at various granularities without the additional expense of multiple model forward passes for the web-scale databases. FF models also generate independent databases which become prohibitively expense to store and switch in between. $\rm Matryoshka~Representation$} enables adaptive retrieval (AR) which alleviates the need to use full-capacity representations, $d=2048$, for all data and downstream tasks. Lastly, all the vector compression techniques [53, 40] used as part of the ANNS pipelines are complimentary to $\rm Matryoshka~Representations$} and can further improve efficiency-vs-accuracy trade-off.
@@ 4.3.1 Adaptive Retrieval
We benchmark $\rm MRL$} in the adaptive retrieval setting (AR) [44]. For a given query image, we obtain a shortlist, $K=200$, of images from the database using a lower-dimensional representation, eg., $D_{s}=16$ followed by reranking with a higher capacity representation, eg., $D_{r}=2048$. In real-world scenarios where top ranking performance is the key objective, measured with mAP@$k$ where k covers a limited yet crucial real-estate, AR provides significant compute and memory gains over single-shot retrieval with representations of fixed dimensionality. Finally, the most expensive part of AR, as with any retrieval pipeline, is the nearest neighbour search for shortlisting. For example even naive re-ranking of 200 images with 2048 dimensions only costs 400 KFLOPs. While we report exact search cost per query for all of the AR, the shortlisting component of the pipeline can be sped-up using ANNS (HNSW). Appendix I has a detailed discussion on compute cost for exact search, memory overhead of HNSW indices and wall-clock times for both implementations. We note that using HNSW with 32 neighbours for shortlisting does not decrease accuracy during retrieval.
@@
Figure 8 showcases the compute-vs-accuracy trade-off for adaptive retrieval using $\rm Matryoshka~Representations$} compared to single-shot using fixed features with ResNet50 on ImageNet-1K. We observe that all AR settings lie above the Pareto frontier of the single-shot retrieval with varying representation sizes. In particular for ImageNet-1K, we show that the AR model with $D_{s}=16$ & $D_{r}=2048$ is as accurate as single-shot retrieval with $d=2048$ while being $\mathbf{\sim 128\times}$ more efficient in theory and $\mathbf{\sim 14\times}$ faster in practice (compared using HNSW on the same hardware). We show similar trends with ImageNet-4K, but note that we require $D_{s}=64$ given the increased difficulty of the dataset. This results in $\sim 32\times$ and $\sim 6\times$ theoretical and in practice speedups respectively. Lastly, while $K=200$ works well for our adaptive retrieval experiments, we ablate over the shortlist size, $k$ in Appendix K.2 to find that the accuracy gains stop after a point further strengthening the use-case for $\rm Matryoshka~Representation~Learning$} and adaptive retrieval.
Even with adaptive retrieval, it is hard to determine the choice of $D_{s}$ & $D_{r}$. In order to alleviate this issue to an extent, we propose Funnel Retrieval, a consistent cascade for adaptive retrieval. Funnel thins out the initial shortlist by a repeated re-ranking and shortlisting with a series of increasing capacity representations. Funnel, halves the shortlist size and double the representation size at every step of re-ranking. For example on ImageNet-1K, a funnel with the shortlist progression of $200\to 100\to 50\to 25\to 10$ with the cascade of $16\to 32\to 64\to 128\to 256\to 2048$ representation sizes within $\rm Matryoshka~Representation$} is as accurate as the single-shot 2048-dim retrieval while being $\sim 128\times$ more efficient theoretically (see Appendix F for more results). All these results showcase the potential of $\rm MRL$} and AR for large-scale multi-stage search systems [15].
@@ 5 Further Analysis and Ablations
We evaluate the robustness of the $\rm MRL$} models trained on ImageNet-1K on out-of-domain datasets, ImageNetV2/R/A/Sketch [64, 32, 33, 85], and compare them to the FF baselines. Table 17 in Appendix H demonstrates that $\rm Matryoshka~Representations$} for classification are at least as robust as the original representation while improving the performance on ImageNet-A by $0.6\%$ – a $20\%$ relative improvement. We also study the robustness in the context of retrieval by using ImageNetV2 as the query set for ImageNet-1K database. Table 9 in Appendix E shows that $\rm MRL$} models have more robust retrieval compared to the FF baselines by having up to $3\%$ higher mAP@$10$ performance. We also find that the zero-shot robustness of ALIGN-$\rm MRL$} (Table 18 in Appendix H) agrees with the observations made by Wortsman et al. [87]. Lastly, Table 6 in Appendix D.2 shows that $\rm MRL$} also improves the cosine similarity span between positive and random image-text pairs.
@@
We exhaustively evaluate few-shot learning on $\rm MRL$} models using nearest class mean [70]. Table 15 in Appendix G shows that that representations learned through $\rm MRL$} perform comparably to FF representations across varying shots and number of classes.
$\rm Matryoshka~Representations$} realize a unique pattern while evaluating on FLUID [83], a long-tail sequential learning framework. We observe that $\rm MRL$} provides up to $2\%$ accuracy higher on novel classes in the tail of the distribution, without sacrificing accuracy on other classes (Table 16 in Appendix G). Additionally we find the accuracy between low-dimensional and high-dimensional representations is marginal for pretrain classes. We hypothesize that the higher-dimensional representations are required to differentiate the classes when few training examples of each are known. This results provides further evidence that different tasks require varying capacity based on their difficulty.
@@
The information packing in $\rm Matryoshka~Representations$} often results in gradual increase of accuracy with increase in capacity. However, we observe that this trend is not ubiquitous and certain instances and classes are more accurate when evaluated with lower-dimensions (Figure 10 in Appendix J). With perfect routing of instances to appropriate dimension, $\rm MRL$} can gain up to $4.6\%$ classification accuracy. At the same time, the low-dimensional models are less accurate either due to confusion within the same superclass [23] of the ImageNet hierarchy or presence of multiple objects of interest. Figure 9 showcases 2 such examples for $8$-dimensional representation. These results along with Appendix J put forward the potential for $\rm MRL$} to be a systematic framework for analyzing the utility and efficiency of information bottlenecks.
@@ 5.1 Ablations
Table 26 in Appendix K presents that $\rm Matryoshka~Representations$} can be enabled within off-the-shelf pretrained models with inexpensive partial finetuning thus paving a way for ubiquitous adoption of $\rm MRL$}. At the same time, Table 27 in Appendix C indicates that with optimal weighting of the nested losses we could improve accuracy of lower-dimensions representations without accuracy loss. Lastly, Tables 28 and 29 in Appendix K.2 show that the retrieval performance saturates after a certain shortlist dimension and length depending on the complexity of the dataset.
Figure 9: Grad-CAM [71] progression of predictions in $\rm MRL$} model across $8,16,32,2048$ dimensions. (a) 88-dimensional representation confuses due to presence of other relevant objects (with a larger field of view) in the scene and predicts “shower cap” & (b) 88-dim model confuses within the same super-class of “boa”; thus failing gracefully in both these scenarios.
@@ 6 Discussion and Conclusions
The results in Section 5.1 reveal interesting weaknesses of $\rm MRL$} that would be logical directions for future work. (1) Optimizing the weightings of the nested losses to obtain a Pareto optimal accuracy-vs-efficiency trade-off. (2) Using different losses at various fidelities aimed at solving a specific aspect of adaptive deployment – e.g. high recall for $8$-dimension and robustness for $2048$-dimension. (3) Learning a search data-structure, like differentiable k-d tree, on top of $\rm Matryoshka~Representation$} to enable dataset and representation aware retrieval. (4) Finally, the joint optimization of multi-objective $\rm MRL$} combined with end-to-end learnable search data-structure to have data-driven adaptive large-scale retrieval for web-scale search applications.
@@ Acknowledgments
We are grateful to Srinadh Bhojanapalli, Lovish Madaan, Raghav Somani and Ludwig Schmidt for helpful discussions and feedback. Part of the paper’s large-scale experimentation is supported through a research GCP credit award from Google Cloud and Google Research. Ali Farhadi acknowledges funding from the NSF awards IIS 1652052, IIS 17303166, DARPA N66001-19-2-4031, DARPA W911NF-15-1-0543 and gifts from Allen Institute for Artificial Intelligence.
@@
      Did you discuss any potential negative societal impacts of your work? [N/A] To be Filled
@@ Appendix C $\rm Matryoshka~Representation~Learning$} Model Training
  • FF-k: ResNet50 model with the fc layer replaced by torch.nn.Linear(k, num_classes), where k $\in[8,16,32,64,128,256,512,1024,2048]$. We will henceforth refer to these models as simply FF, with the k value denoting representation size.
@@ Appendix D Classification Results
We compared the performance of $\rm MRL$} models at various representation sizes via 1-nearest neighbors (1-NN) image classification accuracy on ImageNet-1K in Table 2 and Figure 3. We provide detailed information regarding the k-NN search pipeline in Appendix E. We compared against a baseline of attempting to enforce nesting to a FF-2048 model by 1) Random Feature Selection (Rand. FS): considering the first $m$ dimensions of FF-2048 for NN lookup, and 2) FF+SVD: performing SVD on the FF-2048 representations at the specified representation size. We also compared against the 1-NN accuracy of slimmable neural nets [91] as an additional baseline. We observed these baseline models to perform very poorly at lower dimensions, as they were not explicitly trained to learn $\rm Matryoshka~Representations$}.
@@
Rep. Size $F$ Rand. FS FF+SVD Slimmable $\rm MRL$} $\rm MRL\text{--}E$}
8 58.93  2.36 19.14  1.00 62.19 57.45
16 66.77 12.06 46.02  5.12 67.91 67.05
32 68.84 32.91 60.78 16.95 69.46 68.60
64 69.41 49.91 67.04 35.60 70.17 69.61
128 69.35 60.91 69.63 51.16 70.52 70.12
256 69.72 65.75 70.67 60.61 70.62 70.36
512 70.18 68.77 71.06 65.82 70.82 70.74
1024 70.34 70.41 71.22 67.19 70.89 71.07
2048 71.19 71.19 71.21 66.10 70.97 71.21

### v3

# Matryoshka Representation Learning

@@ 2 Related Work
$\rm MRL$} can be adapted seamlessly to most representation learning frameworks at web-scale with minimal modifications (Section 4.1). For example, $\rm MRL$}’s adaptation to masked language modelling reduces to $\rm MRL\text{--}E$} due to the weight-tying between the input embedding matrix and the linear classifier. For contrastive learning, both in context of vision & vision + language, $\rm MRL$} is applied to both the embeddings that are being contrasted with each other. The presence of normalization on the representation needs to be handled independently for each of the nesting dimension for best results (see Appendix C for more details).
@@ 4.2.1 Adaptive Classification
Figure 7 shows the comparison between cascaded $\rm MRL$} representations ($\rm MRL$}–AC) and independently trained fixed feature (FF) models on ImageNet-1K with ResNet50. We computed the expected representation size for $\rm MRL$}–AC based on the final dimensionality used in the cascade. We observed that $\rm MRL$}–AC was as accurate, $76.30\%$, as a 512-dimensional FF model but required an expected dimensionality of $\sim 37$ while being only $0.8\%$ lower than the 2048-dimensional FF baseline. Note that all $\rm MRL$}–AC models are significantly more accurate than the FF baselines at comparable representation sizes. $\rm MRL$}–AC uses up to $\sim 14\times$ smaller representation size for the same accuracy which affords computational efficiency as the label space grows [84]. Lastly, our results with $\rm MRL$}–AC indicate that instances and classes vary in difficulty which we analyze in Section 5 and Appendix J.
@@ 4.3 Retrieval
$\rm MRL$} models are capable of performing accurate retrieval at various granularities without the additional expense of multiple model forward passes for the web-scale databases. FF models also generate independent databases which become prohibitively expense to store and switch in between. $\rm Matryoshka~Representations$} enable adaptive retrieval (AR) which alleviates the need to use full-capacity representations, $d=2048$, for all data and downstream tasks. Lastly, all the vector compression techniques [57, 43] used as part of the ANNS pipelines are complimentary to $\rm Matryoshka~Representations$} and can further improve the efficiency-vs-accuracy trade-off.
@@ 4.3.1 Adaptive Retrieval
We benchmark $\rm MRL$} in the adaptive retrieval setting (AR) [48]. For a given query image, we obtained a shortlist, $K=200$, of images from the database using a lower-dimensional representation, e.g. $D_{s}=16$ followed by reranking with a higher capacity representation, e.g. $D_{r}=2048$. In real-world scenarios where top ranking performance is the key objective, measured with mAP@$k$ where k covers a limited yet crucial real-estate, AR provides significant compute and memory gains over single-shot retrieval with representations of fixed dimensionality. Finally, the most expensive part of AR, as with any retrieval pipeline, is the nearest neighbour search for shortlisting. For example, even naive re-ranking of 200 images with 2048 dimensions only costs 400 KFLOPs. While we report exact search cost per query for all AR experiments, the shortlisting component of the pipeline can be sped-up using ANNS (HNSW). Appendix I has a detailed discussion on compute cost for exact search, memory overhead of HNSW indices and wall-clock times for both implementations. We note that using HNSW with 32 neighbours for shortlisting does not decrease accuracy during retrieval.
@@
Figure 8 showcases the compute-vs-accuracy trade-off for adaptive retrieval using $\rm Matryoshka~Representations$} compared to single-shot using fixed features with ResNet50 on ImageNet-1K. We observed that all AR settings lied above the Pareto frontier of single-shot retrieval with varying representation sizes. In particular for ImageNet-1K, we show that the AR model with $D_{s}=16$ & $D_{r}=2048$ is as accurate as single-shot retrieval with $d=2048$ while being $\mathbf{\sim 128\times}$ more efficient in theory and $\mathbf{\sim 14\times}$ faster in practice (compared using HNSW on the same hardware). We show similar trends with ImageNet-4K, but note that we require $D_{s}=64$ given the increased difficulty of the dataset. This results in $\sim 32\times$ and $\sim 6\times$ theoretical and in-practice speedups respectively. Lastly, while $K=200$ works well for our adaptive retrieval experiments, we ablated over the shortlist size $k$ in Appendix K.2 and found that the accuracy gains stopped after a point, further strengthening the use-case for $\rm Matryoshka~Representation~Learning$} and adaptive retrieval.
Even with adaptive retrieval, it is hard to determine the choice of $D_{s}$ & $D_{r}$. In order to alleviate this issue to an extent, we propose Funnel Retrieval, a consistent cascade for adaptive retrieval. Funnel thins out the initial shortlist by a repeated re-ranking and shortlisting with a series of increasing capacity representations. Funnel halves the shortlist size and doubles the representation size at every step of re-ranking. For example on ImageNet-1K, a funnel with the shortlist progression of $200\to 100\to 50\to 25\to 10$ with the cascade of $16\to 32\to 64\to 128\to 256\to 2048$ representation sizes within $\rm Matryoshka~Representation$} is as accurate as the single-shot 2048-dim retrieval while being $\sim 128\times$ more efficient theoretically (see Appendix F for more results). All these results showcase the potential of $\rm MRL$} and AR for large-scale multi-stage search systems [15].
@@ 5 Further Analysis and Ablations
We evaluate the robustness of the $\rm MRL$} models trained on ImageNet-1K on out-of-domain datasets, ImageNetV2/R/A/Sketch [68, 32, 33, 89], and compare them to the FF baselines. Table 17 in Appendix H demonstrates that $\rm Matryoshka~Representations$} for classification are at least as robust as the original representation while improving the performance on ImageNet-A by $0.6\%$ – a $20\%$ relative improvement. We also study the robustness in the context of retrieval by using ImageNetV2 as the query set for ImageNet-1K database. Table 9 in Appendix E shows that $\rm MRL$} models have more robust retrieval compared to the FF baselines by having up to $3\%$ higher mAP@$10$ performance. This observation also suggests the need for further investigation into robustness using nearest neighbour based classification and retrieval instead of the standard linear probing setup. We also find that the zero-shot robustness of ALIGN-$\rm MRL$} (Table 18 in Appendix H) agrees with the observations made by Wortsman et al. [91]. Lastly, Table 6 in Appendix D.2 shows that $\rm MRL$} also improves the cosine similarity span between positive and random image-text pairs.
@@
We exhaustively evaluated few-shot learning on $\rm MRL$} models using nearest class mean [74]. Table 15 in Appendix G shows that that representations learned through $\rm MRL$} perform comparably to FF representations across varying shots and number of classes.
$\rm Matryoshka~Representations$} realize a unique pattern while evaluating on FLUID [87], a long-tail sequential learning framework. We observed that $\rm MRL$} provides up to $2\%$ accuracy higher on novel classes in the tail of the distribution, without sacrificing accuracy on other classes (Table 16 in Appendix G). Additionally we find the accuracy between low-dimensional and high-dimensional representations is marginal for pretrain classes. We hypothesize that the higher-dimensional representations are required to differentiate the classes when few training examples of each are known. This results provides further evidence that different tasks require varying capacity based on their difficulty.
Figure 9: Grad-CAM [75] progression of predictions in $\rm MRL$} model across $8,16,32\text{ and }2048$ dimensions. (a) 88-dimensional representation confuses due to presence of other relevant objects (with a larger field of view) in the scene and predicts “shower cap” ; (b) 88-dim model confuses within the same super-class of “boa” ; (c) 88 and $16$-dim models incorrectly focus on the eyes of the doll ("sunglasses") and not the "sweatshirt" which is correctly in focus at higher dimensions; $\rm MRL$} fails gracefully in these scenarios and shows potential use cases of disagreement across dimensions.
@@
The information packing in $\rm Matryoshka~Representations$} often results in gradual increase of accuracy with increase in capacity. However, we observed that this trend was not ubiquitous and certain instances and classes were more accurate when evaluated with lower-dimensions (Figure 12 in Appendix J). With perfect routing of instances to appropriate dimension, $\rm MRL$} can gain up to $4.6\%$ classification accuracy. At the same time, the low-dimensional models are less accurate either due to confusion within the same superclass [23] of the ImageNet hierarchy or presence of multiple objects of interest. Figure 9 showcases 2 such examples for $8$-dimensional representation. These results along with Appendix J put forward the potential for $\rm MRL$} to be a systematic framework for analyzing the utility and efficiency of information bottlenecks.
Figure 10: 31-way ImageNet-1K superclass classification across representation size for $\rm MRL$} & FF models showing the capture of underlying hierarchy through tight information bottlenecks.
Figure 11: Diverse per-superclass accuracy trends across representation sizes for ResNet50-$\rm MRL$} on ImageNet-1K.
Superclass Accuracy.
As the information bottleneck becomes smaller, the overall accuracy on fine-grained classes decreases rapidly (Figure 3). However, the drop-off is not as significant when evaluated at a superclass level (Table 24 in Appendix J). Figure 11 presents that this phenomenon occurs with both $\rm MRL$} and FF models; $\rm MRL$} is more accurate across dimensions. This shows that tight information bottlenecks while not highly accurate for fine-grained classification, do capture required semantic information for coarser classification that could be leveraged for adaptive routing for retrieval and classification. Mutifidelity of $\rm Matryoshka~Representation$} naturally captures the underlying hierarchy of the class labels with one single model. Lastly, Figure 11 showcases the accuracy trends per superclass with $\rm MRL$}. The utility of additional dimensions in distinguishing a class from others within the same superclass is evident for “garment” which has up to 11% improvement for 8 $\to$ 16 dimensional representation transition. We also observed that superclasses such as “oscine (songbird)” had a clear visual distinction between the object and background and thus predictions using 8 dimensions also led to a good inter-class separability within the superclass.
@@ 5.1 Ablations
Table 26 in Appendix K presents that $\rm Matryoshka~Representations$} can be enabled within off-the-shelf pretrained models with inexpensive partial finetuning thus paving a way for ubiquitous adoption of $\rm MRL$}. At the same time, Table 27 in Appendix C indicates that with optimal weighting of the nested losses we could improve accuracy of lower-dimensions representations without accuracy loss. Tables 29 and 29 in Appendix C ablate over the choice of initial granularity and spacing of the granularites. Table 29 reaffirms the design choice to shun extremely low dimensions that have poor classification accuracy as initial granularity for $\rm MRL$} while Table 29 confirms the effectiveness of logarthmic granularity spacing inspired from the behaviour of accuracy saturation across dimensions over uniform. Lastly, Tables 30 and 31 in Appendix K.2 show that the retrieval performance saturates after a certain shortlist dimension and length depending on the complexity of the dataset.
@@ 6 Discussion and Conclusions
The results in Section 5.1 reveal interesting weaknesses of $\rm MRL$} that would be logical directions for future work. (1) Optimizing the weightings of the nested losses to obtain a Pareto optimal accuracy-vs-efficiency trade-off – a potential solution could emerge from adaptive loss balancing aspects of anytime neural networks [39]. (2) Using different losses at various fidelities aimed at solving a specific aspect of adaptive deployment – e.g. high recall for $8$-dimension and robustness for $2048$-dimension. (3) Learning a search data-structure, like differentiable k-d tree, on top of $\rm Matryoshka~Representation$} to enable dataset and representation aware retrieval. (4) Finally, the joint optimization of multi-objective $\rm MRL$} combined with end-to-end learnable search data-structure to have data-driven adaptive large-scale retrieval for web-scale search applications.
@@ Acknowledgments
We are grateful to Srinadh Bhojanapalli, Lovish Madaan, Raghav Somani and Ludwig Schmidt for helpful discussions and feedback. Aditya Kusupati also thanks Tom Duerig and Rahul Sukthankar for their support. Part of the paper’s large-scale experimentation is supported through a research GCP credit award from Google Cloud and Google Research. Gantavya Bhatt is supported in part by the CONIX Research Center, one of six centers in JUMP, a Semiconductor Research Corporation (SRC) program sponsored by DARPA. Sham Kakade acknowledges funding from the NSF award CCF-1703574 and ONR N00014-22-1-2377. Ali Farhadi acknowledges funding from the NSF awards IIS 1652052, IIS 17303166, DARPA N66001-19-2-4031, DARPA W911NF-15-1-0543 and gifts from Allen Institute for Artificial Intelligence.
@@
      Did you discuss any potential negative societal impacts of your work? [N/A] Our work does not have any additional negative societal impact on top of the existing impact of representation learning. However, a study on the trade-off between representation size and the tendency to encode biases is an interesting future direction along the lines of existing literature [34, 35]. A part of this is already presented in Section 5.
@@ Appendix C $\rm Matryoshka~Representation~Learning$} Model Training
  • FF–k: ResNet50 model with the fc layer replaced by torch.nn.Linear(k, num_classes), where k $\in[8,16,32,64,128,256,512,1024,2048]$. We will henceforth refer to these models as simply FF, with the k value denoting representation size.
@@ Appendix D Classification Results
We compared the performance of $\rm MRL$} models at various representation sizes via 1-nearest neighbors (1-NN) image classification accuracy on ImageNet-1K in Table 2 and Figure 3. We provide detailed information regarding the k-NN search pipeline in Appendix E. We compared against a baseline of attempting to enforce nesting to a FF-2048 model by 1) Random Feature Selection (Rand. FS): considering the first $m$ dimensions of FF-2048 for NN lookup, and 2) FF+SVD: performing SVD on the FF-2048 representations at the specified representation size, 3) FF+JL: performing random projection according to the Johnson-Lindenstrauss lemma [46] on the FF-2048 representations at the specified representation size. We also compared against the 1-NN accuracy of slimmable neural nets [95] as an additional baseline. We observed these baseline models to perform very poorly at lower dimensions, as they were not explicitly trained to learn $\rm Matryoshka~Representations$}.
@@
Rep. Size Rand. FS SVD JL $F$ Slimmable $\rm MRL$} $\rm MRL\text{--}E$}
8  2.36 19.14 0.11 58.93  1.00 62.19 57.45
16 12.06 46.02 0.09 66.77  5.12 67.91 67.05
32 32.91 60.78 0.06 68.84 16.95 69.46 68.6
64 49.91 67.04 0.05 69.41 35.60 70.17 69.61
128 60.91 69.63 0.06 69.35 51.16 70.52 70.12
256 65.75 70.67 0.04 69.72 60.61 70.62 70.36
512 68.77 71.06 0.03 70.18 65.82 70.82 70.74
1024 70.41 71.22 - 70.34 67.19 70.89 71.07
2048 71.19 71.21 - 71.19 66.10 70.97 71.21